OpenStudy (anonymous):

cos(theta)-tan(theta)cos(thetha)=0

OpenStudy (anonymous):

How do you find the values of this problem?

OpenStudy (anonymous):

\[\tan(\Theta) = \frac{ \sin(\Theta) }{ \cos(\Theta) }\] just substitute that back for tan(theta) and simplify

OpenStudy (anonymous):

ok i got cos(theta)-(sin(theta)/cos(theta))cos(theta)=0 cos(theta)-sin(theta)=0 now how do i find the values, cause the answers are a. 0, pi/4, pi, 5pi/4 b.pi/4, 5pi/4 c.pi/2, 3pi/4, 3pi/2, 7pi/4 d. pi/2, 7pi/6, 3pi/2, 11pi/6

OpenStudy (anonymous):

You know what I should have done.... factor the cos(theta) so this is easier to solve \[\cos(\Theta) [ 1 - \tan(\Theta)] = 0\]

OpenStudy (anonymous):

so now solve for \[\cos(\Theta) = 0\] and \[\tan(\Theta) = -1 \]

OpenStudy (anonymous):

so it would be c for my answer from the above choices listed?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i submitted the answer and it was wrong, the correct answer was b but i dont understand how that is, because choice c actually makes sense.

OpenStudy (anonymous):

Oh no! I missed a - sign it should have been \[1- \tan(\Theta) = \tan(\theta) = 1 \]

OpenStudy (anonymous):

but how is the answer b.. they completely ignored the cos(theta)?

OpenStudy (anonymous):

i have no idea, im going to ask my counselor if they could reset it or maybe if it was a mistake or something, any ways thankyou for your help